Calculate the weight of an astronaut whose mass (including spacesuit) is 72 kg on the

Moon. Mass of the Moon = 7.4 × 1022 kg. Radius of Moon = 1.74 × 106 m.


1
Expert's answer
2021-07-05T15:23:06-0400

The weight of an astronaut 

"P=mg=G\frac{mM}{R^2}"

"P=6.67\times 10^{-11}\times\frac{72\times 7.4\times 10^{22}}{(1.74\times 10^6)^2}=117\:\rm N"